A 250 g block slides on a rough horizontal table. Find the work done by the frictional force in bringing the block to rest if it is initially moving at a speed of 40 cm/s. If the friction coefficient between the table and the block is 0.1, how far does the block move before coming to rest?

Grade:11

1 Answers

Deepak Patra
9 years ago
Sol. . Given, m = 25 base o g = 0.250kg, u = 40 cm/sec = 0.4m/sec μ = 0.1, v=0 Here, μ R = ma {where, a = deceleration} a = μR/m = μmg/m = μg = 0.1 * 9.8 = 0.98 m/sec^2 S = v^2 – u^2/2a = 0.082m = 8.2 cm Again, work done against friction is given by - W = μ RS cos θ = 0.1 * 2.5 * 0.082 * 1 (θ = 0°) = 0.02 J ⇒ W = - 0.02 J